OpenStudy (anonymous):

5 = 76 + BLANK Correct answer gets medal and fan

OpenStudy (anonymous):

pretty sure its a negative

OpenStudy (texaschic101):

5 = 76 + x --- subtract 76 from both sides 5 - 76 = x - 71 = x check.. 5 = 76 - 71 5 = 5 (correct) answer is -71

OpenStudy (anonymous):

wow